### v1

# Large Language Models as Automatic Annotators and Annotation Adjudicators for Fine-Grained Opinion Analysis

Abstract
Fine-grained opinion analysis of text provides a detailed understanding of expressed sentiments, including the addressed entity. Although this level of detail is sound, it requires considerable human effort and substantial cost to annotate opinions in datasets for training models, especially across diverse domains and real-world applications. We explore the feasibility of LLMs as automatic annotators for fine-grained opinion analysis, addressing the shortage of domain-specific labelled datasets. In this work, we use a declarative annotation pipeline. This approach reduces the variability of manual prompt engineering when using LLMs to identify fine-grained opinion spans in text. We also present a novel methodology for an LLM to adjudicate multiple labels and produce final annotations. After trialling the pipeline with models of different sizes for the Aspect Sentiment Triplet Extraction (ASTE) and Aspect-Category-Opinion-Sentiment (ACOS) analysis tasks, we show that LLMs can serve as automatic annotators and adjudicators, achieving high Inter-Annotator Agreement across individual LLM-based annotators. This reduces the cost and human effort needed to create these fine-grained opinion-annotated datasets.
1 Introduction
There has been a considerable evolution in the field of fine-grained opinion analysis, which has led to many theoretical and practical formulations of opinions and sentiment. These fine-grained methods provide a clearer picture of not just the sentiment polarity (positive, negative, neutral) but also the contextual elements of the aforementioned sentiments. Consider the following example:
“I had hoped for better battery life, as it had only about 2-1/2 hours doing heavy computations (8 threads using 100% of the CPU)”
In the text, a negative sentiment is being expressed. However, a fine-grain annotation, unlike sentence or document-level annotation, has the opinion span expressing the polarity, i.e. “hoped for better”, along with an explicit instantiation of the target in the text called an aspect term “battery life”, forming an aspect sentiment triplet as specified according to the Aspect Sentiment Triplet Extraction (ASTE) task [16]. Moreover, the annotation can also include a more coarse-grained property “Battery#Operational_performance”, which is an aspect category, and expresses the attribute of the entity towards which the opinion is specifically directed. Aspect categories enable the encoding of cases where the aspect term is missing and the target sentiment is expressed implicitly. In addition to triplets elements, the quadruple annotations are assigned categories too, adhering to Aspect-Category-Opinion-Sentiment (ACOS) guidelines [1].
The existing datasets used for fine-grained opinion mining tasks are derived from SemEval tasks and are relatively small, limited to specific domains, and collected exclusively from review websites. As highlighted in [18], these datasets are helpful in comparing different ABSA methods but are not sufficient for addressing real-world use cases. The subjective annotation of opinions by human annotators is a non-trivial task and requires strict adherence to annotation guidelines, as documented in [13].
LLMs are few-shot learners [8] and exhibit instruction following and reasoning capabilities [2], which warrants the assessment of their use for annotating opinions. In this work, we investigate the use of LLMs to annotate text with fine-grained opinions. We also propose a novel LLM-based Adjudication approach that combines multiple opinion annotations for a given text input to resolve inter-annotator disagreement and obtain final opinion annotations. Our main contributions include:
  • Evaluation of fine-grained opinion annotation accuracy of LLMs (on ASTE and ACOS)
  • Declarative LLM-driven opinion annotation pipeline using DSPy.
  • We propose a method for aggregating annotations from multiple LLM-driven artificial annotators to produce final annotations using LLM as an annotation adjudicator.
2 Related Work
2.1 Opinion Formulations
Opinion [^1] Unless stated otherwise, we use the term opinion as a broad concept that covers sentiment and its associated information, such as opinion target and the person who holds the opinion, and use the term sentiment to mean only the underlying positive, negative, or neutral polarity implied by opinion. Mining and sentiment analysis have been well explored in natural language processing. Aspect-Based Sentiment Analysis (ABSA), the foundation of fine-grained opinion mining, evolved from feature-based summarisation [4, 10], which involves extracting and summarising opinions about features (attributes/ keywords). In this paper, we experiment with two fine-grained and extractive formulations of downstream ABSA tasks: Aspect Sentiment Triple Extraction (ASTE) [17, 16] and Aspect-Category-Opinion-Sentiment Quadruple (ACOS) extraction [1].
2.2 LLMs as Annotators
Key steps in dataset creation for building any intelligent system include data selection, guideline development, human annotation, quality estimation, and adjudication [7]. With the improvement of the general ability of LLMs, they have been investigated for annotating various subjective tasks, including span annotations[5], argument quality annotations[11] and propaganda span annotations[3]. Their findings argue in favour of using LLMs as independent or additional annotators.
2.3 Prompt Engineering
LLMs are known to be sensitive to the prompts [19] and exhibit a higher level of variability in performing complex tasks [14]. To minimise the spurious interaction between the prompt and model selection as a confounding variable, we adopt DSPy [6] and the approach of “programming LLMs” rather than prompting them naively. It shifts the focus from performing string manipulation on prompt strings to programming with structured, declarative modules that automatically prepare prompts using a small sample of annotated examples.
In this work, we propose a declarative pipeline based on LLMs for fine-grained opinion annotation. In addition to utilising individual LLMs to annotate text with fine-grained opinions, we also propose a novel LLM-based Adjudication approach to combine multiple annotations to resolve inter-annotator disagreement and obtain final annotations.
3 Data
3.1 Data Models
We use two fine-grained formulations to describe data models for opinion representation (Figure 1(a)); these are described subsequently.
ASTE
Aspect Sentiment Triple Extraction (ASTE) falls under the umbrella of ABSA tasks. Proposed by [12], it is a task for extracting opinion triplets, thus providing one-shot answer to: What target is being discussed, How is the sentiment (e.g. positive) and Why is this sentiment (e.g. “excellent”).
ACOS
This data model is based on named Aspect-Category-Opinion-Sentiment (ACOS) Quadruple Extraction, to extract aspect, aspect category, opinion and sentiment as quadruples in text and provide full support for aspect-based sentiment analysis with implicit aspects and opinions  [^2] Opinion in ACOS task refers to the span that expresses sentiment.. The primary focus of this formulation is the fine-grained analysis of the opinion target. This task is more challenging than ASTE because it not only adds a category on top of the existing opinion triplet, but also takes into account implicit cases (i.e. missing explicit aspect terms or opinion spans) [1].
(a) ACOS & ASTE Specifications
Task Dataset Train Dev Test
ASTE lap14 906 219 328
res14 1126 310 492
res15 607 148 322
res16 857 210 326
ACOS laptop 2934 326 816
restaurant 1530 171 583
(b) Dataset breakdown
3.2 Datasets
To align our experiments with those of the original ASTE and ACOS tasks, we use subsections of the original dataset splits, as shown in Table 1(b). We disregard the training splits for each dataset as we are not fine-tuning or training the LLMs in our pipeline. For the development splits, we use random sampling to select half of the entries as possible in-context learning (ICL) examples required by the optimal prompt, then use the other half of the development set for evaluating which number of ICLs (5, 10, or 15) in the prompt produces the best performing model. This division is equal across all datasets’ development sets, except for ASTE’s lap14 and ACOS’s restaurant datasets, which each have one more example in their prompt evaluation than in their prompt creation. Finally, the test splits are kept the same to evaluate performance across multiple LLMs in our experiments.
4 Method
The fine-grained opinion annotation pipeline has multiple stages: (i) Preparing and evaluating LLM-specific prompts, (ii) Redundant Opinion annotation using multiple LLMs, and (iii) Annotation adjudication using LLM to create collaborative opinion annotations.
4.1 Problem Statements
The two steps of annotating opinions using the fine-grained opinion formulations involve identifying a set of opinions expressed in the input text $T_{i}$. As with human annotators, the process generates multiple annotations for the same input $T_{i}$, followed by adjudication that resolves disagreements between annotators to produce final annotations.
LLM as opinion annotators Given the input text $Ti$ from corpora $\mathcal{T}$, our objective is to extract all expressed opinions $Oi=\{o_{1},o_{2},..o_{n}\}$. The elements of single opinion ($o_{j}$) is task dependant:
  • ASTE opinions are formulated as the triple $o_{j}=\{at,s,op\}$, where $at$ is the aspect term, $s$ is the sentiment and $op$ is the opinion term.
  • ACOS opinions are quadruples $o_{j}=\{at,s,op,ac\}$, where $at$, $s$ and $op$ encode similar concepts as ASTE and $ac$ is aspect category made up of entity $E$ and attribute $A$ (expressed as E#AE\#A).
The LLM-driven annotation pipeline maps input corpora $\mathcal{T}$ to $\mathcal{O}^{k}$ using $k$ LLM-based annotators.
LLM as annotation adjudicators We use LLM to perform the adjudication to combine opinions annotated by $k$ different LLMs to obtain the final opinion annotation, i.e., $\mathcal{A}\mathrm{dj}_{\mathrm{LLM}}:(\mathcal{T},\mathcal{O}^{1},\dots,\mathcal{O}^{k})\rightarrow\mathcal{O}$. This method is motivated by an ensemble machine learning technique called stacking [15].
4.2 Declarative Annotation Pipeline
Figure 2: LLM-Based Annotation Pipeline using DSPy (Left) with LLM-as-adjudicator (Right)
The proposed annotation pipeline as illustrated in Fig 2 uses LLMs for automatic annotation. However, LLMs are known to be sensitive to the prompts [19] and exhibit a higher level of variability in performing complex tasks [14]. To minimise the spurious interaction between the prompt and model selection as a confounding variable, we adopt the DSPy [6] method of programming LLMs rather than prompting them naively. DSPy’s pipeline outputs an optimised prompt and the mechanism to facilitate data input and generated output in our workflow.
Prior to annotating the dataset, we use DSPy to generate an optimal prompt. It uses a data model to provide a general description of its inputs (text) and expected output, along with the data types. We use a small number of annotated examples from the Dev set for this purpose, giving one half to the DSPy program as a pool of ICL candidates to construct optimal prompts for the LLM using 5, 10, or 15 annotated examples from the Dev set. Next, we use the other half of the Dev set to evaluate each prompt and determine the number of ICL examples that is most effective for the model. Once the final configuration has been selected, the optimal DSPy model is run on the test set to produce a set of opinion annotations. It should be noted that in our pipeline, we never fine-tune the LLM. We repeat this across multiple LLMs across multiple sizes to obtain multiple annotations for the test data. Finally, we repeat this process to optimise an adjudication prompt that resolves redundant annotations into the final annotations.
5 Experimental Setup
In our experiments, we used models of three different sizes: mini ( 4B parameters), small ( 14B parameters), and medium ( 32B parameters). The assignment of LLMs as $\mathcal{A}_{1},\mathcal{A}_{1},\mathcal{A}_{3}$ is carried out based on the model’s performance on evaluation samples and for all cases, the adjudication is performed by $\mathcal{A}_{1}$, which is the notation assigned to the best-performing LLM-based annotator.
Annotator Mini ( 4B) Small ( 14B) Medium ( 32B)
$\mathcal{A}_{1}$ Qwen3-4B-Thinking Qwen3-14B Qwen/Qwen3-30B-A3B-Thinking-2507-FP8
$\mathcal{A}_{2}$ MiniCPM3-4B openai/gpt-oss-20b unsloth/DeepSeek-R1-Distill-Qwen-32B-bnb-4bit
$\mathcal{A}_{3}$ Phi-4-mini-reasoning DeepSeek-R1-Distill-Qwen-14B openai/gpt-oss-20b
Table 1: LLMs used as annotators
Each experiment was run on a single Nvidia A40 GPU, with vLLM [9] used to serve each model through the various steps of our pipeline, as required by DSPy for OpenAI-compliant API access. We keep the LLM hyperparameters constant across all models. The temperature is set to 0.0 to ensure the most deterministic generation. The model’s context window determines the input sequence length, and the generated sequence’s output length is set to 16384. Since the objective of the experiments is to conduct annotations using LLMs’ common-sense understanding, we do not use the training data. We use half of the small evaluation set provided for DSPy prompt optimisation to generate prompts with 5, 10, and 15 ICL examples, then use the remaining evaluation set examples to select the best-performing DSPy-wrapped prompt for annotating the test split.
We evaluate the performance of multiple LLM-based annotation models on the ASTE and ACOS tasks using standard exact match metrics. For each dataset, we report precision (P), recall (R), and F1-score across different data splits or subsets, labelled as Mini, Small, and Med.
6 Results
The trends observed in the LLM-based annotations, compared with the human annotations, are discussed individually for both annotation tasks. The evaluation presented in this section reports the triplet and quadruplet F1-scores for ASTE and ACOS, respectively. These scores are computed using human annotations as a reference and reflect their alignment with LLM-based annotations.
lap14 res14 res15 res16
$\mathcal{A}_{1}$ $\mathcal{A}_{2}$ $\mathcal{A}_{3}$ $\mathcal{A}_{\mathrm{dj}}$ $\mathcal{A}_{1}$ $\mathcal{A}_{2}$ $\mathcal{A}_{3}$ $\mathcal{A}_{\mathrm{dj}}$ $\mathcal{A}_{1}$ $\mathcal{A}_{2}$ $\mathcal{A}_{3}$ $\mathcal{A}_{\mathrm{dj}}$ $\mathcal{A}_{1}$ $\mathcal{A}_{2}$ $\mathcal{A}_{3}$ $\mathcal{A}_{\mathrm{dj}}$
Mini P 42.34 34.05 21.63 45.95 62.27 54.71 36.18 67.44 50.96 41.70 30.12 54.31 62.43 52.42 40.65 64.28
R 43.39 38.08 25.92 45.43 64.98 60.16 41.22 67.85 60.60 55.46 40.87 62.65 70.81 67.51 51.27 71.92
F1 43.15 35.95 23.59 45.69 63.86 57.31 38.53 67.64 55.11 47.61 34.68 58.18 66.36 59.01 45.34 67.89
Small P 44.05 40.07 31.79 38.89 62.39 57.00 55.26 59.13 54.92 47.49 45.88 50.66 60.33 53.20 57.66 57.01
R 45.85 42.14 34.20 41.77 64.49 63.48 59.76 64.38 66.80 62.47 56.29 63.56 70.43 66.15 68.09 69.64
F1 44.93 41.08 32.95 40.24 63.42 60.06 57.42 61.65 60.28 53.96 50.56 56.38 64.99 58.98 62.44 62.69
Med P 43.93 42.31 40.07 45.14 66.38 60.38 57.00 68.81 56.77 55.03 47.49 57.66 59.87 58.25 53.20 62.33
R 47.50 44.73 42.14 47.50 71.12 64.69 63.48 70.42 68.67 66.60 62.47 68.24 71.98 68.68 66.15 72.76
F1 45.64 43.49 41.08 46.81 68.67 62.46 50.06 69.27 61.98 60.26 53.96 62.51 65.37 63.04 58.98 67.14
Table 2: LLM-based annotator performances on ASTE (triplet exact match)
ASTE: As reported in Table 2, Individual LLM-based opinion annotators generally achieve better alignment with human annotations as their parameter size increases, a trend that is strong across most annotators. The adjudication process further improves this alignment for Mini and Med models. While individual Mini models show variable performance, their adjudicated results can surpass even non-adjudicated Small and Med models. For Small models, individual annotators converge more closely with human annotations than  4B Mini-sized models do. Still, adjudication surprisingly reduces alignment, likely due to the inclusion of the non-"thinking" Qwen3-14B model. Ultimately, the Medium-sized (referred to as Med) models achieved the best individual alignment, which was then improved through adjudication to yield the overall best results across all datasets.
Model Laptop Restaurant
Mini Small Med Mini Small Med
P R F1 P R F1 P R F1 P R F1 P R F1 P R F1
$\mathcal{A}_{1}$ 13.18 12.97 13.07 9.78 9.60 9.69 15.77 15.74 15.75 32.11 28.71 30.31 22.67 20.19 21.36 29.37 27.51 28.41
$\mathcal{A}_{2}$ 11.65 12.63 12.12 7.62 7.69 7.66 9.63 9.94 9.78 21.27 21.51 21.39 24.27 21.72 22.92 17.48 17.46 17.47
$\mathcal{A}_{3}$ 1.36 1.90 1.58 6.92 6.49 6.70 7.62 7.69 7.66 4.62 5.89 5.18 16.88 14.74 15.73 24.27 21.72 22.92
$\mathcal{A}\mathrm{dj}$ 16.71 16.17 16.43 10.94 11.07 11.01 18.31 19.72 18.99 32.90 30.45 31.63 30.86 29.14 29.98 41.12 39.95 40.53
Table 3: LLM-based annotator performances on ACOS (quadruple exact match)
ACOS: Table 3 shows that, as with ASTE, the disparity amongst LLM annotators decreases as the number of parameters in the models increases and the Med models produce the annotations most aligned with the human annotated labels. Adjudication consistently improved results across all model scales, unlike in the ASTE task.
The Mini $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ models show promising performance when compared to the Small scale, however the lowest performing annotations overall are by Mini $\mathcal{A}_{3}$.
The lower overall scores highlight the greater challenge of annotating quadruples, a task further complicated by domain differences: the laptop domain, with 114 distinct categories $ac$, proved more difficult than the restaurant domain, with only 12, adversely affecting the exact-match F1 score on the full quadruples.
7 Analysis
Data $\mathcal{A}$ Mini Small Med
$s$ atat opop s&ats\&at s&ops\&op at&opat\&op $s$ atat opop s&ats\&at s&ops\&op at&opat\&op $s$ atat opop s&ats\&at s&ops\&op at&opat\&op
lap14 $\mathcal{A}_{1}$ 83.55 66.74 64.02 58.22 56.53 48.67 85.80 66.18 65.93 59.56 60.36 48.91 84.81 66.12 72.73 58.02 64.64 51.33
$\mathcal{A}_{2}$ 80.81 63.21 50.98 53.82 45.61 39.82 83.29 64.30 61.88 55.96 53.97 47.03 85.80 64.46 68.75 56.49 60.52 49.96
$\mathcal{A}_{3}$ 79.77 51.34 37.39 43.30 32.01 26.99 86.25 58.32 49.31 52.3 44.12 36.87 83.29 64.30 61.88 55.96 53.97 47.03
Adj 83.46 67.38 69.24 57.62 61.55 52.62 84.21 63.81 60.43 56.07 53.82 45.63 84.44 67.36 72.91 58.03 64.37 53.37
res14 $\mathcal{A}_{1}$ 90.86 79.93 75.50 75.64 71.77 67.13 89.94 78.35 77.00 73.95 73.06 66.43 90.44 81.35 79.38 77.53 75.73 71.10
$\mathcal{A}_{2}$ 85.61 77.74 68.55 70.72 62.91 61.69 88.21 77.08 72.02 72.48 67.76 62.42 89.53 78.65 75.79 73.73 71.70 65.86
$\mathcal{A}_{3}$ 86.49 63.91 53.05 59.00 49.63 40.38 90.50 77.81 68.51 73.69 65.24 59.86 88.19 77.35 72.29 72.75 67.97 62.83
Adj 90.63 80.47 79.97 75.97 76.05 70.59 90.24 78.56 73.58 74.11 69.96 64.78 90.20 81.64 79.82 77.28 75.98 72.24
res15 $\mathcal{A}_{1}$ 88.49 76.80 72.87 69.92 64.91 61.19 90.53 82.39 73.86 75.62 67.52 66.11 89.50 77.44 78.25 71.00 71.30 67.79
$\mathcal{A}_{2}$ 83.26 75.46 63.20 66.19 54.44 54.37 88.71 78.43 68.71 71.80 61.82 59.70 89.47 78.92 76.62 72.42 69.38 66.04
$\mathcal{A}_{3}$ 85.96 60.81 51.07 53.82 46.01 38.76 91.01 79.43 65.75 73.85 59.06 55.76 88.74 78.52 68.83 71.92 61.96 59.84
Adj 88.48 79.02 76.14 71.10 68.37 65.13 89.85 80.56 70.76 73.62 63.94 62.63 89.69 79.82 78.27 73.73 71.63 67.99
res16 $\mathcal{A}_{1}$ 88.75 81.83 78.93 76.14 73.16 71.23 89.79 79.35 79.61 73.74 74.29 69.48 88.98 78.83 80.55 73.16 74.68 70.14
$\mathcal{A}_{2}$ 88.83 75.00 72.00 69.78 67.05 62.69 87.14 76.36 75.34 70.04 69.56 63.49 90.91 76.98 78.16 72.18 73.35 66.79
$\mathcal{A}_{3}$ 87.91 65.34 57.80 61.07 53.83 47.58 89.97 79.01 74.42 74.11 70.54 65.83 87.14 76.36 75.34 70.04 69.56 63.49
Adj 89.55 81.96 81.33 76.29 75.70 72.49 89.04 78.09 77.21 72.44 71.77 67.08 89.73 80.04 82.44 74.62 76.45 72.17
Table 4: Element-wise assessment of ASTE’s alignment with human annotations
The results in Table 3 and 2 joint evaluation (i.e. triplets for ASTE and quadruples for ACOS). In this section we investigate the individual labels forming triplets and quadruples, and identify which combinations are less performant.
Dataset Annot $s$ atat opop $E$ AA acac s&ats\&at s&ops\&op at&opat\&op s&acs\&ac at&acat\&ac op&acop\&ac at&s&opat\&s\&op
Mini laptop $\mathcal{A}_{1}$ 89.48 68.97 57.05 53.05 54.05 30.75 63.47 54.96 42.50 28.26 21.68 18.06 40.65
$\mathcal{A}_{2}$ 88.52 62.30 45.09 66.95 54.12 34.99 56.61 42.59 30.39 31.43 23.54 18.19 28.55
$\mathcal{A}_{3}$ 73.39 37.60 17.16 46.63 39.58 19.53 32.25 15.98 10.42 16.43 7.24 3.40 9.91
Adj 89.70 68.45 55.21 67.81 54.83 39.87 63.12 52.30 40.37 36.29 27.76 22.75 38.50
Restaurant $\mathcal{A}_{1}$ 90.34 68.69 65.58 62.51 76.59 55.87 63.70 61.79 51.20 51.85 40.53 39.00 48.75
$\mathcal{A}_{2}$ 87.77 61.60 55.47 57.80 65.34 49.39 58.54 52.76 41.05 45.86 31.35 29.70 39.51
$\mathcal{A}_{3}$ 76.86 38.22 23.84 34.68 49.03 26.51 34.30 21.89 15.26 23.54 13.85 8.46 14.27
Adj 90.48 69.26 58.19 85.88 75.92 66.25 64.77 54.08 45.64 61.83 50.14 40.86 43.00
Small laptop $\mathcal{A}_{1}$ 90.88 59.26 36.86 65.58 59.28 40.56 55.24 35.22 22.85 37.48 23.66 16.83 21.88
$\mathcal{A}_{2}$ 89.30 57.66 44.10 49.61 49.31 21.97 53.06 41.23 27.21 19.92 15.54 10.62 25.65
$\mathcal{A}_{3}$ 90.62 63.35 38.24 48.28 50.50 21.86 58.98 35.99 25.16 20.13 13.67 10.19 23.89
Adj 90.85 60.04 34.75 76.28 60.71 46.74 55.77 32.77 22.59 43.32 27.08 18.10 21.37
restaurant $\mathcal{A}_{1}$ 90.45 57.85 46.60 67.11 78.10 58.95 54.22 44.18 34.11 54.98 37.52 27.80 32.59
$\mathcal{A}_{2}$ 87.93 60.00 49.27 54.07 68.22 49.11 56.03 46.46 37.32 44.55 36.23 27.91 35.69
$\mathcal{A}_{3}$ 90.26 64.57 50.73 51.98 67.26 43.78 60.42 46.77 36.58 39.51 28.45 22.93 34.34
Adj 90.08 64.18 47.57 79.63 84.32 71.48 59.87 44.72 37.71 66.00 49.19 37.46 35.78
Med laptop $\mathcal{A}_{1}$ 89.86 68.20 55.29 62.96 53.05 36.30 62.67 52.17 39.57 33.35 25.95 22.61 37.46
$\mathcal{A}_{2}$ 90.37 61.84 43.96 56.42 56.75 33.15 57.22 42.07 29.04 30.59 20.72 15.40 27.66
$\mathcal{A}_{3}$ 89.30 57.66 44.10 49.61 49.31 21.97 53.06 41.23 27.21 19.92 15.54 10.62 25.65
Adj 90.40 73.53 59.59 70.51 54.27 41.23 67.71 56.60 45.80 37.90 30.68 24.84 43.53
restaurant $\mathcal{A}_{1}$ 89.94 69.39 64.13 65.41 72.28 54.91 63.90 60.16 50.09 49.01 39.26 37.66 47.46
$\mathcal{A}_{2}$ 91.50 63.30 54.71 52.97 63.72 46.25 58.98 52.52 39.54 42.48 28.08 25.09 38.32
$\mathcal{A}_{3}$ 87.93 60.00 49.27 54.07 68.22 49.11 56.03 46.46 37.32 44.55 36.23 27.91 35.69
Adj 91.01 72.74 64.60 82.96 82.43 71.75 68.13 61.17 52.86 66.40 55.78 48.85 50.06
Table 5: Element-wise assessment of ACOS alignment with human annotations (using F1 score)
In Table 4, sentiment polarity ($s$) predictions have the highest alignment with respect to human annotations. There are fewer $at$ and $op$ spans than $s$, but they are very comparable in alignment. The AT extraction and assignment of correct polarity (s&ats\&at) is more challenging than individually labelling $s$ and $at$, but it is less challenging than $s$ and $op$ which is expected as sentiment polarity and words used to express the polarity are easier to predict jointly. Based on these scores, the most challenging step short of predicting the entire triplet jointly is at&opat\&op identification (i.e. correctly identifying which $op$ span expresses the $op$ on which target.
Similar trends to the ASTE triplets can be found in Table 5. All the models short of 32B exhibit high domain disparity, which can be clearly attributed to the previously discussed difficulty of correctly predicting $ac$.
7.1 Inter-Annotator Agreement of LLM-Based Annotators
In accordance with standard annotation practices, we report the pre-adjudication Inter-Annotator Agreement (IAA) to evaluate LLM-based annotators’ reliability. While F1-scores are used for the human-annotated dataset, their limitations as an IAA metric are noted in [7]. Thus, we employ Krippendorff’s $\alpha$ scores to assess the IAAs of our LLM-based span annotation.
Model lap14 res14 res15 res16
atat opop at&sat\&s op&sop\&s atat opop at&sat\&s op&sop\&s atat opop at&sat\&s op&sop\&s atat opop at&sat\&s op&sop\&s
Mini 0.63 0.49 0.59 0.46 0.70 0.61 0.68 0.59 0.70 0.57 0.67 0.54 0.72 0.63 0.69 0.60
Small 0.73 0.66 0.72 0.66 0.79 0.74 0.78 0.74 0.81 0.72 0.80 0.71 0.79 0.77 0.78 0.75
Med 0.75 0.69 0.75 0.69 0.82 0.77 0.80 0.77 0.84 0.75 0.82 0.74 0.82 0.79 0.81 0.78
Model Laptop Restaurant
atat opop acac at&sat\&s op&sop\&s atat opop acac at&sat\&s op&sop\&s
Mini 0.60 0.37 0.22 0.58 36.75 0.61 0.41 0.41 0.59 0.40
Small 0.70 0.65 0.31 0.69 0.65 0.73 0.64 0.49 0.71 0.64
Med 0.70 0.64 0.33 0.69 0.64 0.71 0.65 0.50 0.69 0.64
Table 6: Top: Krippendorff-$\alpha$ for IAA for ASTE span annotations (atat: Aspect, opop: Opinion, at&sat\&s: Aspect Term with sentiment, op&sop\&s: Opinion span with sentiment). Bottom: Performance scores for ACOS tasks (acac: Aspect Category).
Upon examining the Krippendorff $\alpha$ values, we observe that the reliability of the labels for ASTE and ACOS increases as the model parameter size increases. For ASTE, we see the reliability of triplet span elements approaching highly reliable levels, whereas ACOS quadruple elements tend to be moderately reliable at best. The aforementioned trend highlights the observed increase in annotation reliability with increasing model size.
7.2 Error Analysis
In the Table 7, we observe how the entire prediction in ASTE is deemed incorrect in triplet evaluation due to a minor mistake in identifying the aspect term ($ac$). In the ACOS formulation for the same example, we see the challenge model faces in identifying $at$.
In the second example, while ASTE is correctly predicted, we see challenges in identifying implicit values: the second quadruple fails to recognise that the aspect term is implicit, instead predicting an implicit opinion span. This is what makes ACOS more challenging than the ASTE task.
Text ASTE Pred (Med ADJ) ASTE Gold ACOS Pred (Med ADJ) ACOS Gold
I have eaten here three times and have found the quality and variety of the fish to be excellent. $s$: Positive $at$: quality $op$: excellent $s$: Positive $at$: fish $op$: excellent $s$: Positive $at$: quality $ac$: food#quality $op$: excellent $s$: Positive $at$: variety $ac$: food#style $op$: excellent $s$: Positive $at$: fish $ac$: food#quality $op$: excellent $s$: Positive $at$: fish $ac$: food#style $op$: excellent
All their menu items are a hit, and they serve mimosas. $s$: Positive $at$: menu items $op$: hit $s$: Positive $at$: menu items $op$: hit $s$: Positive $at$: menu items $ac$: food#quality $op$: hit $s$: Neutral $at$: mimosas $ac$: drinks#general $op$: null $s$: Positive $at$: menu items $ac$: food#quality $op$: hit $s$: Positive $at$: null $ac$: drinks#style $op$: hit
Table 7: Error Analysis: Opinions on identical text based on ASTE, ACOS formulation along with the predictions
8 Conclusion
This work demonstrates that Large Language Models can effectively serve as automatic annotators, reducing the high cost and effort of creating human-annotated datasets for fine-grained opinion analysis. By implementing a declarative DSPy pipeline for the annotation process and introducing a novel LLM-based adjudication method, we have shown that LLMs can produce high-quality, reliable annotations for ASTE opinion triplets and ACOS quadruples. We evaluated the human alignment of the produced annotations using the F1-score and found that ACOS tasks show greater deviation from human annotations than ASTE tasks do. We further investigate and discuss the challenges that make ACOS more difficult than ASTE. As customary with human annotators, we report the Kirppendorff $\alpha$ to assess the reliability of the annotated dataset. Our results across various model sizes confirm the viability of this approach, paving the way for a more scalable, cost-effective method to build detailed datasets by aiding human annotators or providing additional annotators.
Acknowledgment
This research work is funded by Research Ireland under Grant Number SFI/12/RC/2289_P2 (Insight), co-funded by the European Regional Development Fund.